# Learning to Drive Ethically: Embedding Moral Reasoning into Autonomous Driving

### v1

@@ Abstract
Autonomous vehicles hold great promise for reducing traffic fatalities and improving transportation efficiency, yet their widespread adoption hinges on embedding robust ethical reasoning into routine and emergency maneuvers. Here, we present a hierarchical Safe Reinforcement Learning (Safe RL) framework that explicitly integrates moral considerations with standard driving objectives. At the decision level, a Safe RL agent is trained using a composite ethical risk cost, combining collision probability and harm severity, to generate high-level motion targets. A dynamic Prioritized Experience Replay mechanism amplifies learning from rare but critical, high-risk events. At the execution level, polynomial path planning coupled with Proportional–Integral–Derivative (PID) and Stanley controllers translates these targets into smooth, feasible trajectories, ensuring both accuracy and comfort. We train and validate our approach on rich, real-world traffic datasets encompassing diverse vehicles, cyclists, and pedestrians, and demonstrate that it outperforms baseline methods in reducing ethical risk and maintaining driving performance. To our knowledge, this is the first study of ethical decision-making for autonomous vehicles via Safe RL in real-world scenarios. Our results highlight the potential of combining formal control theory and data-driven learning to advance ethically accountable autonomy in complex, human-mixed traffic environments.
@@ 1 Introduction
Over the past several decades, both academic institutions and industry stakeholders have devoted significant effort to advancing autonomous driving (AD) technologies. By eliminating human error, autonomous vehicles (AVs) promise to reduce traffic accidents and enhance both social welfare and economic productivity. While fully functional Level 5 systems, with completely autonomous vehicles available on the mass market, are expected to emerge in the coming decades (Khan et al., 2022), the transitional period will involve extensive interaction between AVs and human-driven vehicles. This raises a pressing societal question: Can AVs make responsible decisions in emergencies, especially when interacting with human drivers (Othman, 2021)? Beyond emergency responses, broader ethical and policy concerns have also emerged, including issues of legal liability, data privacy, and equitable access to mobility solutions (Wang et al., 2020a; Lin, 2016; Bonnefon et al., 2016). Early public and academic debates often focused on abstract dilemmas such as the “trolley problem”, how an AV should react when a collision is unavoidable, empirical work such as the Moral Machine Experiment has demonstrated that ethical preferences vary widely across cultures, for example, favoring younger over older pedestrians or humans over animals (Awad et al., 2018). However, this emphasis on rare, extreme scenarios risks obscuring the everyday ethical decisions AVs must make, such as maintaining safe following distances, navigating busy intersections, and yielding to pedestrians, which cumulatively determine public safety and perceived fairness in urban environments (Evans et al., 2020). Moreover, special attention must be paid to protecting vulnerable road users, whose risk of severe injury remains disproportionately high in the event of a collision (Campolettano et al., 2024).
@@
To that end, we propose EthicAR, an ethical AD agent built upon the principles of Safe RL. EthicAR separates the objectives of AV into two components: standard driving goals (e.g., safety and efficiency) and moral reasoning across surrounding traffic participants. This structure allows the vehicle to make ethically informed decisions, particularly in the face of potential collisions. The framework is organized into two hierarchical levels: a decision level and an execution level. The decision level determines how and where the vehicle should move, producing a target lateral position and target longitudinal speed. It is implemented using a Safe RL algorithm that incorporates ethical risk as a cost signal, following the formulation introduced in (Geisslinger et al., 2023). The risk metric combines two factors: (i) the probability of collision involving either the ego vehicle or other road users, and (ii) the severity of harm associated with those collisions. Because high-risk situations are rare in naturalistic driving data, we introduce a dynamic Prioritized Experience Replay (PER) mechanism. This strategy emphasizes learning from high-risk experiences, thereby enhancing the ability to learn from rare but critical situations. The execution level translates high-level decisions into physical motion. It includes a path planner that generates smooth, feasible trajectories using polynomial equations. A low-level path follower then executes these trajectories using a PID controller for longitudinal speed and a Stanley controller for lateral position tracking. This setup ensures stable and comfortable vehicle behavior. To train the agent, we use real-world datasets that capture diverse traffic scenarios with various types of vehicles, pedestrians, and cyclists. These datasets provide realistic interactions, allowing the learned policy to generalize effectively to heterogeneous and dynamic environments. We evaluate EthicAR against several baselines in real-world scenarios and present a comprehensive analysis of its performance. To the best of our knowledge, this is the first study to apply Safe RL to ethical decision-making in AVs using real-world data, backed by extensive empirical validation. In summary, the main contributions of this paper are as follows:
@@ 2 Background
RL is a foundational component of artificial intelligence, where an agent learns to make decisions through trial-and-error interactions with its environment (Sutton and Barto, 2018). This learning process is typically formalized as a Markov Decision Process (MDP), represented by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ denotes the state space, capturing all possible configurations of the environment. $\mathcal{A}$ denotes the action space, describing the set of actions available to the agent. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ models the likelihood of transitioning from one state to another, given an action. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scalar feedback to each state-action pair. The discount factor $\gamma\in[0,1)$ determines the relative importance of future rewards compared to immediate ones. Within this framework, the goal of the agent is to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to a probability distribution over actions, so as to maximize the expected cumulative return:
@@
To assess the effectiveness of a policy $\pi$, two value functions are commonly used. The state-value function, $V^{\pi}(s_{t})=\mathbb{E}\pi\left[R_{t}\mid s_{t}\right]$, estimates the expected return when starting in state $s_{t}$ and following policy $\pi$ thereafter. The action-value function, $Q^{\pi}(s_{t},a_{t})=\mathbb{E}\pi\left[R_{t}\mid s_{t},a_{t}\right]$, measures the expected return of taking action $a_{t}$ in state $s_{t}$, then following $\pi$. Based on how these value functions are utilized, RL methods are broadly categorized into two groups: value-based and policy gradient methods. In value-based approaches, an optimal policy $\pi^{\star}$ is derived by acting greedily to the optimal action-value function $Q^{\star}$. This optimal function satisfies the condition $Q^{\star}(s,a)=\max_{\pi}Q^{\pi}(s,a)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Conversely, policy gradient methods aim to directly optimize a parameterized policy $\pi_{\theta}$ by maximizing a performance objective, such as the expected return under $\pi_{\theta}$. A prominent subclass of these methods is the actor-critic architecture, where the actor represents the policy and the critic estimates value functions to guide the updates of the actor (Haarnoja et al., 2018).
@@
where $\pi_{\theta}$ is a parameterized policy. To solve such constrained problems, Lagrangian relaxation (Boyd and Vandenberghe, 2004) and projection-based policy updates (Achiam et al., 2017) are commonly used. The Lagrangian method introduces a non-negative dual variable $\lambda\geq 0$ to convert the constrained problem into an unconstrained saddle-point formulation:
@@ 3.1 Problem Definition
In this paper, we propose an AD framework that leverages RL to generate safe and socially compliant trajectories. The framework continuously processes real-time environmental data, including interactions with other vehicles, cyclists, and pedestrians. Its goal is to balance efficiency, safety, and ethical behavior. To achieve this, we design a two-level control architecture with a decision level and an execution level. The decision level uses Safe RL to generate high-level motion targets. The execution level translates these targets into smooth and comfortable vehicle control.
The RL agent operates within a Frenet coordinate system, outputting a policy $\pi$, that maps observed states $s_{t}\in\mathcal{S}$ to actions $a_{t}=(T_{t},D_{d,t},v_{s,t})\in\mathcal{A}$, where $T_{t}$ denotes planning time, $D_{d,t}$ defines lateral displacement, and $v_{s,t}$ specifies target longitudinal speed. The observation and action space, detailed in Section 4.2, ensures compatibility with dynamic road geometries. Using these parameters, the planner generates collision-free trajectories $\xi=\left\{(s_{i},d_{i})\right\}_{i=1}^{N}$, which the follower executes via low-level control (Section 4.3).
A key contribution of our framework is the integration of ethical risk minimization into the RL objective. Unlike traditional RL approaches, which focus solely on performance or safety, our method explicitly aims to reduce risks to all road users. Building on Safe RL principles introduced earlier, we formulate this as a constrained optimization problem. The agent seeks to maximize the expected return $\mathcal{J}_{\pi}$ while minimizing a cumulative ethical risk metric, defined as $D_{ethics}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}_{ethics}(s_{t},a_{t},s_{t+1})\right]$, where $\mathcal{C}_{\text{ethics}}$ quantifies potential risk to pedestrians, cyclists, and other road users. By incorporating this ethical cost, described in detail in Section 3.5, directly into the constraint, our approach ensures that planned trajectories reflect not only a commitment to collision avoidance but also a broader sense of moral responsibility. For example, the agent learns to reduce unnecessary proximity to vulnerable individuals or to distribute unavoidable risk more equitably. This ethical integration sets our method apart from conventional AV planning techniques, which often treat safety and ethics as secondary considerations.
@@ 3.2 LSTM based SACLag
Soft Actor-Critic (SAC) algorithm combines actor-critic architecture with entropy regularization. It aims to maximize both cumulative rewards and policy entropy, encouraging exploration by favoring stochastic policies. SAC employs a temperature parameter $\alpha$ to balance reward maximization and entropy trade-offs. Crucially, this temperature is automatically adjusted to maintain a target entropy level, avoiding manual tuning. The actor optimizes a stochastic policy, while the critic estimates Q-values using clipped double-Q networks to reduce overestimation bias (Haarnoja et al., 2018). SAC with Lagrangian relaxation (SACLag) uses a Lagrangian multiplier to optimize while enforcing the cost constraint dynamically. It transforms the problem into a dual optimization: the primal variables (policy and Q-networks) maximize the reward-entropy objective, while the dual variable (Lagrangian multiplier) is adjusted via gradient ascent on the Lagrange loss.
As outlined in Section 3.1, the objective of the RL agent is to dynamically supply the path planner with relevant information for trajectory generation. However, individual states frequently lack the contextual depth needed to support optimal decision-making by the agent. To address this limitation, we integrate Long Short-Term Memory (LSTM) networks into the SACLag framework. Each network in this framework incorporates LSTM layers (Hochreiter and Schmidhuber, 1997), which process sequences of past time steps. This temporal processing capability allows the agent to retain and leverage historical information alongside the current state, thereby enabling more informed and consistent decisions.
@@ 3.3 Collision probabilities
To estimate the potential risk to other road users, we begin by computing the collision probabilities between the ego vehicle and its surrounding road users. Since our framework performs trajectory planning over a fixed time horizon, we evaluate these probabilities at each timestep within that horizon. Given the planned trajectory of the ego vehicle, the next step is to predict the future trajectories of nearby road users. For surrounding vehicles, we use an LSTM-based neural network that incorporates convolutional social pooling and semantic road context, using bird’s-eye-view images as input (Geisslinger et al., 2021). For additional implementation details, we refer readers to (Geisslinger et al., 2021). For other types of road users, such as pedestrians and cyclists, we use their ground-truth trajectories with added Gaussian noise to account for uncertainty. The predicted trajectories of these agents are modeled as Gaussian distributions, denoted as $\xi_{o}=[\mathbf{\mu}_{o,\tau},\mathbf{\Sigma}_{o,\tau}]_{\tau=1}^{T}$, with $\mathbf{\mu}_{\tau},\mathbf{\Sigma}_{\tau}]$ represent the position mean and covariance matrix at time $\tau$.
@@
For the first stage, together with the planned trajectories, denoted as $\xi_{e}=[\textbf{p}_{e,\tau}]_{\tau=1}^{T}$, from our framework, we then compute overlap probabilities $P_{sat}$ using SAT. The SAT states that two convex shapes do not intersect if at least one axis exists along which the projections of the shapes are completely separated; conversely, if no such axis exists, the shapes must be colliding. This theorem is operationalized by determining potential separating axes, typically the normals to the edges of the convex bodies, in our context, oriented bounding box (OBB), then projecting each shape onto these axes to assess the intervals of projection; if every projection interval overlaps, a collision is confirmed, while the existence of even one non-overlapping projection interval serves as evidence of separation. This method has been widely adopted in the design of physics engines and simulation frameworks due to its computational efficiency and robustness in handling various orientations and configurations of convex objects. Let $\mathcal{A}^{\tau}=\left\{a_{1}^{\tau},...,a_{m}^{\tau}\right\}$ be the set of separating axes for the OBB pair at time $\tau$. The overlap condition is:
$\mathcal{O}_{e}^{\tau}\cap\mathcal{O}_{o}^{\tau}\neq\emptyset\Longleftrightarrow\forall a\in\mathcal{A}^{\tau}:\text{proj}_{a}(\mathcal{O}_{e}^{\tau})\cap\text{proj}_{a}(\mathcal{O}_{o}^{\tau})\neq\emptyset,$ (4)
@@
$P_{sat}^{\tau}=\mathbb{E}\left\{\mathbbm{1}\left(\bigcap_{a\in\mathcal{A}^{\tau}}[\text{proj}_{a}(\mathcal{O}_{e}^{\tau})\cap\text{proj}_{a}(\mathcal{O}_{o}^{\tau})\neq\emptyset]\right)\right\},$ (5)
@@
If an overlap is detected, we compute the Mahalanobis distance between the ego vehicle and the predicted distribution of the position of other traffic participants. The distance is a measure of how far a point is from a distribution, taking into account the shape of the distribution. Assuming collisions at different timesteps are independent, the Mahalanobis distance between two traffic participants at each time $\tau$ is then expressed as:
@@
The squared Mahalanobis distance follows a chi-squared distribution with 2 degrees of freedom, i.e., $(D_{m,\tau})^{2}\sim\chi^{2}(2)$. We convert the distance into a collision probability using the complementary cumulative distribution function of the chi-squared distribution:

### v2

# Learning to Drive Ethically: Moral Reasoning in Autonomous Driving

@@ Abstract
Autonomous vehicles hold great promise for reducing traffic fatalities and improving transportation efficiency, yet their widespread adoption hinges on embedding robust ethical reasoning into routine and emergency maneuvers, particularly to protect vulnerable road users (VRUs) such as pedestrians and cyclists. Here, we present a hierarchical Safe Reinforcement Learning (Safe RL) framework that explicitly integrates moral considerations with standard driving objectives. At the decision level, a Safe RL agent is trained using a composite ethical risk cost, combining collision probability and harm severity, to generate high-level motion targets. A dynamic Prioritized Experience Replay mechanism amplifies learning from rare but critical, high-risk events. At the execution level, polynomial path planning coupled with Proportional–Integral–Derivative (PID) and Stanley controllers translates these targets into smooth, feasible trajectories, ensuring both accuracy and comfort. We train and validate our approach on rich, real-world traffic datasets encompassing diverse vehicles, cyclists, and pedestrians, and demonstrate that it outperforms baseline methods in reducing ethical risk and maintaining driving performance. To our knowledge, this is the first study of ethical decision-making for autonomous vehicles via Safe RL evaluated on real-world, human-mixed traffic scenarios. Our results highlight the potential of combining formal control theory and data-driven learning to advance ethically accountable autonomy that explicitly protects those most at risk in urban traffic environments.
@@ 1 Introduction
Over the past several decades, both academic institutions and industry stakeholders have devoted significant effort to advancing autonomous driving (AD) technologies. By eliminating human error, autonomous vehicles (AVs) promise to reduce traffic accidents and enhance both social welfare and economic productivity. While fully functional Level 5 systems, with completely autonomous vehicles available on the mass market, are expected to emerge in the coming decades (Khan et al., 2022), the transitional period will involve extensive interaction between AVs and human-driven vehicles. This raises a pressing societal question: Can AVs make responsible decisions in emergencies and everyday interactions, not only to protect their occupants but also to protect those outside the vehicle (Othman, 2021)? This question is especially urgent for vulnerable road users (VRUs) such as pedestrians and cyclists, who bear a disproportionately high burden of severe injury and death in collisions. According to the U.S. National Highway Traffic Safety Administration, 7,522 pedestrians were killed and more than 67,000 injured in 2022 (Naumann, 2025). Safeguarding VRUs should therefore be a primary objective of technical design, regulation, and public policy (Campolettano et al., 2024). Beyond emergency responses, broader ethical and policy concerns have also emerged, including legal liability, data privacy, and equitable access to mobility solutions (Wang et al., 2020a; Lin, 2016; Bonnefon et al., 2016). Early public and academic debates often focused on abstract dilemmas such as the “trolley problem”, how an AV should react when a collision is unavoidable. Empirical work such as the Moral Machine Experiment has demonstrated that ethical preferences vary widely across cultures, for example, favoring younger over older pedestrians or humans over animals (Awad et al., 2018). However, emphasizing these rare, extreme scenarios risks obscuring the everyday ethical decisions AVs must make, such as maintaining safe following distances, navigating busy intersections, and yielding to pedestrians, which cumulatively determine public safety and perceived fairness in urban environments (Evans et al., 2020).
@@
To that end, we propose EthicAR, an ethical AD agent built upon the principles of Safe RL. EthicAR separates the objectives of AV into two components: standard driving goals (e.g., safety and efficiency) and moral reasoning across surrounding traffic participants. This structure allows the vehicle to make ethically informed decisions, particularly in the face of potential collisions. The framework is organized into two hierarchical levels: a decision level and an execution level. The decision level does not directly control the vehicle, instead, it determines how and where the vehicle should move, producing a target lateral position and target longitudinal speed. It is implemented using a Safe RL algorithm that incorporates ethical risk as a cost signal, following the formulation introduced in (Geisslinger et al., 2023). The risk metric combines two factors: (i) the probability of collision involving either the ego vehicle or other road users, and (ii) the severity of harm associated with those collisions. Because high-risk situations are rare in naturalistic driving data, we introduce a dynamic Prioritized Experience Replay (PER) mechanism. This strategy emphasizes learning from high-risk experiences, thereby enhancing the ability to learn from rare but critical situations. The execution level translates high-level decisions into physical motion. It includes a path planner that generates smooth, feasible trajectories using polynomial equations. A low-level path follower then executes these trajectories using a PID controller for longitudinal speed and a Stanley controller for lateral position tracking (Thrun et al., 2006). This setup ensures stable and comfortable vehicle behavior. To train the agent, we use real-world datasets that capture diverse traffic scenarios with various types of vehicles, pedestrians, and cyclists. These datasets provide realistic interactions, allowing the learned policy to generalize effectively to heterogeneous and dynamic environments. We evaluate EthicAR against several baselines in real-world scenarios and present a comprehensive analysis of its performance. To the best of our knowledge, this is the first study to apply Safe RL to ethical decision-making in AVs using real-world data, backed by extensive empirical validation. In summary, the main contributions of this paper are as follows:
@@ 2 Background
RL is a foundational component of artificial intelligence, where an agent learns to make decisions through trial-and-error interactions with its environment (Sutton and Barto, 2018). This learning process is typically formalized as a Markov Decision Process (MDP), represented by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ denotes the state space, capturing all possible configurations of the environment. $\mathcal{A}$ denotes the action space, describing the set of actions available to the agent. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ models the likelihood of transitioning from one state to another, given an action. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scalar feedback to each state-action pair. The discount factor $\gamma\in[0,1)$ determines the relative importance of future rewards compared to immediate ones. Within this framework, the goal of the agent is to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to a probability distribution over actions, so as to maximize the expected cumulative discounted return:
@@
To assess the effectiveness of a policy $\pi$, two value functions are commonly used. The state-value function, $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t}\right]$, estimates the expected return when starting in state $s_{t}$ and following policy $\pi$ thereafter. The action-value function, $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t},a_{t}\right]$, measures the expected return of taking action $a_{t}$ in state $s_{t}$, then following $\pi$. Based on how these value functions are utilized, RL methods are broadly categorized into two groups: value-based and policy gradient methods. In value-based approaches, an optimal policy $\pi^{\star}$ is derived by acting greedily to the optimal action-value function $Q^{\star}$. This optimal function satisfies the condition $Q^{\star}(s,a)=\max_{\pi}Q^{\pi}(s,a)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Conversely, policy gradient methods aim to directly optimize a parameterized policy $\pi_{\theta}$ by maximizing a performance objective, such as the expected return under $\pi_{\theta}$. A prominent subclass of these methods is the actor-critic architecture, where the actor represents the policy and the critic estimates value functions to guide the updates of the actor (Haarnoja et al., 2018).
@@
where $\pi_{\theta}$ is a parametrized policy. To solve such constrained problems, Lagrangian relaxation (Boyd and Vandenberghe, 2004) and projection-based policy updates (Achiam et al., 2017) are commonly used. The Lagrangian method introduces a non-negative dual variable $\lambda\geq 0$ to convert the constrained problem into an unconstrained saddle-point formulation:
@@ 3.1 Problem Definition
In this paper, we propose an AD framework that leverages RL to generate safe and socially compliant trajectories. The framework continuously processes real-time environmental data, including interactions with other vehicles, cyclists, and pedestrians, with more details in Section 4.2. Its goal is to balance efficiency, safety, and ethical behavior. To achieve this, we design a two-level control architecture with a decision level and an execution level. The decision level uses Safe RL to generate high-level motion targets. The execution level translates these targets into smooth and comfortable vehicle control.
The RL agent operates within a Frenet coordinate system, which describes motion relative to a reference path using two parameters: l, the longitudinal distance along the path, and d, the lateral offset from the path. This representation simplifies trajectory planning on curved roads by decoupling progress and deviation, providing an intuitive view of the vehicle’s relationship to the road. The RL agent should output a policy $\pi$, that maps observed states $s_{t}\in\mathcal{S}$ to actions $a_{t}=(T_{t},D_{d,t},v_{l,t})\in\mathcal{A}$, where $T_{t}$ denotes target planning time, $D_{d,t}$ defines target lateral displacement, and $v_{l,t}$ specifies target longitudinal speed along the reference path. The observation and action space, detailed in Section 4.2, are designed to ensure compatibility with dynamic road geometries. Using these outputs, the planner generates collision-free trajectories
$\xi_{e}=[\textbf{p}_{e,\tau}]_{\tau=1}^{T},\quad\textbf{p}_{e,\tau}=[l_{e,\tau},d_{e,\tau}],$
where $\textbf{p}_{e,\tau}$ represents the target position in the Frenet coordinate system of the ego vehicle at each timestep $\tau$ within the target planning time. These trajectories are executed by the follower via low-level control, as described in Section 4.3.
A key contribution of our framework is the integration of ethical risk minimization into the RL objective. Unlike traditional RL approaches, which focus solely on performance or safety, our method explicitly aims to reduce risks to all road users. Building on Safe RL principles introduced earlier, we formulate this as a constrained optimization problem. The agent seeks to maximize the expected return $\mathcal{J}_{\pi}$ while minimizing a cumulative ethical risk metric, defined as $D_{\text{ethics}}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}_{\text{ethics}}(s_{t},a_{t},s_{t+1})\right]$, where $\mathcal{C}_{\text{ethics}}$ quantifies potential risk to pedestrians, cyclists, and other road users. By incorporating this ethical cost, described in detail in Section 3.5, directly into the constraint, our approach ensures that planned trajectories reflect not only a commitment to collision avoidance but also a broader sense of moral responsibility. For example, the agent learns to reduce unnecessary proximity to vulnerable individuals or to distribute unavoidable risk more equitably. This ethical integration sets our method apart from conventional AV planning techniques, which often treat safety and ethics as secondary considerations.
@@ 3.2 LSTM based SACLag
Soft Actor-Critic (SAC) algorithm combines actor-critic architecture with entropy regularization. It aims to maximize both cumulative rewards and policy entropy, encouraging exploration by favoring stochastic policies. SAC employs a temperature parameter $\alpha$ to balance reward maximization and entropy trade-offs. This temperature is automatically adjusted to maintain a target entropy level, avoiding manual tuning. The actor optimizes a stochastic policy, while the critic estimates Q-values using clipped double-Q networks to reduce overestimation bias (Haarnoja et al., 2018). SAC with Lagrangian relaxation (SACLag) uses a Lagrangian multiplier to optimize while enforcing the cost constraint dynamically. It transforms the problem into a dual optimization: the primal variables (policy and Q-networks) maximize the reward-entropy objective, while the dual variable (Lagrangian multiplier) is adjusted via gradient ascent on the Lagrange loss.
As outlined in Section 3.1, the objective of the RL agent is to dynamically supply the path planner with relevant information for trajectory generation. However, individual states frequently lack the contextual depth needed to support optimal decision-making by the agent. In particular, historical information about the ego vehicle and surrounding traffic is important. To formally account for this, we model the problem as a partially observable Markov decision process (POMDP) (Kaelbling et al., 1998), which extends the MDP tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ with an observation space $\mathcal{O}$ and an observation function $\mathcal{Z}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$, where $\Delta(\mathcal{O})$ denotes the probability distribution over $\mathcal{O}$. In a POMDP, the agent does not directly observe the true state $s_{t}\in\mathcal{S}$ but instead receives an observation $o_{t}\in\mathcal{O}$ generated according to $\mathcal{Z}(o_{t}\mid s_{t},a_{t-1})$. To address this partial observability, we integrate Long Short-Term Memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) into the SACLag framework to deal with POMDPs. Each network in this framework incorporates LSTM layers, which process sequences of past time steps. This temporal reasoning capability enables the agent to retain and exploit historical information in combination with the current observation, thereby supporting more informed and consistent decision-making.
@@ 3.3 Collision probabilities
To estimate the potential risk to other road users, we begin by computing the collision probabilities between the ego vehicle and its surrounding road users. Since the framework optimizes trajectories over varied time horizons, these probabilities are evaluated at each discrete timestep within that horizon. Given the planned trajectory of the ego vehicle, the next step is to predict the future trajectories of nearby road users. For surrounding vehicles, we use an LSTM-based neural network that incorporates convolutional social pooling and semantic road context, using bird’s-eye-view images as input (Geisslinger et al., 2021). It provides at each timestep $\tau$ a mean position $\mu_{o,\tau}$ and a covariance matrix $\mathbf{\Sigma}_{o,\tau}$, defined over the longitudinal distance along the path and the lateral displacement. Under the Gaussian assumption, these parameters provide the probability distribution of the predicted positions of the agent. Additional implementation details are provided in (Geisslinger et al., 2021). For other types of agents, such as pedestrians and cyclists, ground-truth trajectories are used with Gaussian noise added to capture uncertainty. The predicted trajectory of agent $o$ is thus represented as a sequence of Gaussian variables,
$\xi_{o,\tau}\sim\mathcal{N}(\mathbf{\mu}_{o,\tau},\mathbf{\Sigma}_{o,\tau}),\quad\tau=0,\dots,T,\quad\xi_{o}=\{\xi_{o,\tau}\}_{\tau=0}^{T}.$
@@
In the first stage, given the planned ego trajectories $\xi_{e}$ from our framework, we compute the overlap probabilities $P_{\text{sat}}$ using the SAT. The SAT states that two convex shapes do not intersect if there exists at least one axis along which their projections are completely separated; conversely, if no such axis exists, the shapes must be colliding. This theorem is operationalized by determining potential separating axes, typically the normals of the edges of the convex bodies. In our context, these convex bodies are oriented bounding boxes (OBBs). Each shape is projected onto the candidate axes, and the intervals of projection are compared: a collision is confirmed if all projection intervals overlap, whereas the existence of even a single non-overlapping interval implies separation. This method has been widely adopted in the design of physics engines and simulation frameworks due to its computational efficiency and robustness in handling various orientations and configurations of convex objects. Let $\mathcal{U}^{\tau}=\left\{u_{1}^{\tau},...,u_{m}^{\tau}\right\}$ be the set of separating axes for the OBB pair at time $\tau$. The overlap condition is:
$\mathcal{O}_{e}^{\tau}\cap\mathcal{O}_{o}^{\tau}\neq\emptyset\Longleftrightarrow\forall u\in\mathcal{U}^{\tau}:\text{proj}_{u}(\mathcal{O}_{e}^{\tau})\cap\text{proj}_{u}(\mathcal{O}_{o}^{\tau})\neq\emptyset,$ (4)
@@
$P_{sat}^{\tau}=\mathbbm{1}\left\{\bigcap_{u\in\mathcal{U}^{\tau}}[\text{proj}_{u}(\mathcal{O}_{e}^{\tau})\cap\text{proj}_{u}(\mathcal{O}_{o}^{\tau})\neq\emptyset]\right\},$ (5)
@@
If an overlap is detected, we compute the Mahalanobis distance between the ego vehicle and the predicted expected position of other traffic participants, taking into account uncertainties in the prediction. Assuming collisions at different timesteps are independent, the Mahalanobis distance between two traffic participants at each time $\tau$ is then expressed as:
@@
Under the assumption of Gaussianity, the squared Mahalanobis distance follows a chi-squared distribution with 2 degrees of freedom, i.e., $(D_{m,\tau})^{2}\sim\chi^{2}(2)$. We convert the distance into a collision probability using the survival distribution function of the chi-squared distribution: